Schedule of Cash Payments

Tadpole Learning Systems Inc. was organized on February 28. Projected selling and administrative expenses for each of the first three months of operations are as follows:

March $133,100
April 126,400
May 115,000

Depreciation, insurance, and property taxes represent $29,000 of the estimated monthly expenses. The annual insurance premium was paid on February 28, and property taxes for the year will be paid in November. 75% of the remainder of the expenses are expected to be paid in the month in which they are incurred, with the balance to be paid in the following month.

Prepare a schedule indicating cash payments for selling and administrative expenses for March, April, and May. Enter all amounts as positive numbers.

TADPOLE LEARNING SYSTEMS INC.
Schedule of Cash Payments for Selling and Administrative Expenses
For the Three Months Ending May 31
March April May
March expenses:
Paid in March $
Paid in April $
April expenses:
Paid in April
Paid in May $
May expenses:
Paid in May
Total cash payments $ $ $
Answer #1

Prepare a schedule indicating cash payments for selling and administrative expenses for March, April, and May. Enter all amounts as positive numbers.

TADPOLE LEARNING SYSTEMS INC.
Schedule of Cash Payments for Selling and Administrative Expenses
For the Three Months Ending May 31
March April May
March expenses:
Paid in March (133100-29000)*75% = 78075
Paid in April $26025
April expenses:
Paid in April (126400-29000)*75% = 73050
Paid in May $24350
May expenses:
Paid in May (115000-29000)*75% = 64500
Total cash payments $78075 $99075 $88850